OpenStudy (anonymous):

How do I write this equation in standard form? (Conic sections) x^2 + y^2 - 8x + 10y + 15 = 0

OpenStudy (anonymous):

Also, I assume identifying the important features would be the center point, foci, etc.?

OpenStudy (anonymous):

i don't know this i'm sorry :(

OpenStudy (anonymous):

that's fine lol

OpenStudy (potatoes.ramu):

Hey! Try grouping the x, and y terms together to start with :)

OpenStudy (anonymous):

Like, grouping them together how? like x^2 - 8x and y^2 + 10y?

OpenStudy (potatoes.ramu):

Yup! exactly! Then you'll get an equation that you can compare to the standard form!

OpenStudy (anonymous):

I'm not quite sure what the standard form is... I'm looking through my notes because I was sure I wrote it down but apparently not lol

OpenStudy (anonymous):

it is a circle

OpenStudy (anonymous):

it has a center and a radius, no foci etc

OpenStudy (anonymous):

you find it by turning \[x^2 + y^2 - 8x + 10y + 15 = 0\] in to something that looks like \[(x-h)^2+(y-k)^2=r^2\] by completing the square twice do you know how to do that?

OpenStudy (anonymous):

I've done it before, completing the square, but I think I could use a refresher

OpenStudy (jdoe0001):

@TammisaurusRex do yo know what a "perfect square trinomial" is?

OpenStudy (anonymous):

It's an equation in the form ax^2 + bx + c or something like that, right?

OpenStudy (jdoe0001):

well.. ahemm ---> ax^2 + bx + c <--- is just a quadratic equation, a 2nd degree equation

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=xGOQYTo9AKY <---